OpenStudy (harleygirl18):

Pretty please help with a cherry on top. Will medal For homework, Fred has to read 30 minutes daily. It takes him 2 minutes per page, and he spends 4 minutes viewing graphics in the book. If x represents the number of pages he reads, then 2x + 4 ≥ 30 represents the time he reads each day. What is the fewest number of pages Fred must read to complete his homework? _____

OpenStudy (harleygirl18):

@surjithayer

OpenStudy (anonymous):

\[2x \ge30-4,2x \ge26,x \ge \frac{ 26 }{ 2 },x \ge 13\] minimum number of pages=?